\smallskip

\begin{thebibliography}{88}

\bibitem{banik-bharali:2024}
Annapurna Banik and Gautam Bharali, {\em Unbounded visibility domains:
metric estimates and an application}. To appear in Ann. Scuola Norm. Super. Pisa Cl. Sci.,
arXiv reference:
{\href{https://doi.org/10.48550/arXiv.2405.05704}{\texttt{arXiv:2405.05704v3}}}.

\bibitem{baracco-khanh-pinton-zampieri:2016}
Luca Baracco, Tran Vu Khanh, Stefano Pinton, and Giuseppe Zampieri,
{\em H{\"o}lder regularity of the solution to the complex Monge--Amp{\`e}re equation with $L^p$ density},
Calc. Var. Partial Differential Equations {\bf 55} (2016), no.~4, Art.~74, 8 pp.


\bibitem{bedford-taylor:1976}
Eric Bedford and B.A. Taylor, {\em The Dirichlet problem for a complex
Monge--Amp{\`e}re equation}, Invent. Math. {\bf 37} (1976), no.~1, 1-44.

\bibitem{bharali-masanta:2025}
Gautam Bharali and Rumpa Masanta, {\em On some connections between Kobayashi geometry and
pluripotential theory}, preprint, arXiv reference:
{\href{https://arxiv.org/abs/2505.16949v2}{\texttt{arXiv:2505.16949v2}}}.

\bibitem{blocki:1996}
Zbigniew B{\l}ocki, {\em The complex Monge--Amp{\`e}re operator in hyperconvex domains},
Ann. Scuola Norm. Sup. Pisa Cl. Sci. {\bf 23} (1996), no.~4, 721-747.

\bibitem{caffarelli-kohn-nirenberg-spruck:1985}
L. Caffarelli, J.J. Kohn, L. Nirenberg, and J. Spruck, {\em The Dirichlet problem for
nonlinear second-order elliptic equations II: Complex Monge--Amp{\`e}re, and uniformly
elliptic, equations}, Comm. Pure Appl. Math. {\bf 38} (1985), no.~2, 209-252.

\bibitem{caffarelli-nirenberg-spruck:1985}
L. Caffarelli, L. Nirenberg, and J. Spruck, {\em The Dirichlet problem for nonlinear
second-order elliptic equations III: Functions of the eigenvalues of the Hessian},\
Acta Math. {\bf 155} (1985), no.~3-4, 261-301.

\bibitem{charabati:2015}
Mohamad Charabati, {\em H{\"o}lder regularity for solutions to complex
Monge--Amp{\`e}re equations}, Ann. Polon. Math. {\bf 113} (2015), no.~2, 109-127.

%\bibitem{charabati:2017}

\bibitem{coman:1997}
Dan Coman, {\em Domains of finite type and H{\"o}lder continuity of the Perron--Bremermann function},
Proc. Amer. Math. Soc. {\bf 125} (1997), no.~12, 3569-3574.

\bibitem{DeVore-Lorentz:1993}
Ronald A. DeVore and George G. Lorentz, Constructive approximation,
Grundlehren der mathematischen Wissenschaften [Fundamental Principles
of Mathematical Sciences], {\bf 303}, Springer-Verlag, Berlin, 1993. 

\bibitem{fornaess-sibony:1989}
John Erik Forn{\ae}ss and Nessim Sibony, {\em Construction of P.S.H. functions on
weakly pseudoconvex domains} Duke Math. J. {\bf 58} (1989), no.~3, 633-655.

\bibitem{guedj-kolodziej-zeriahi:2008}
V. Guedj, S. Ko{\l}odziej, and A. Zeriahi, {\em H{\"o}lder continuous solutions to Monge--Amp{\`e}re
equations}, Bull. Lond. Math. Soc. {\bf 40} (2008), no.~6, 1070-1080.

\bibitem{guedj-zeriahi:2017}
Vincent Guedj and Ahmed Zeriahi, Degenerate complex Monge--Amp{\`e}re equations,
EMS Tracts in Mathematics, {\bf 26}. European Mathematical Society (EMS), Z{\"u}rich, 2017.

\bibitem{ha-khanh:2015}
Ly Kim Ha and Tran Vu Khanh, {\em Boundary regularity of the solution to the
complex Monge--Ampère equation on pseudoconvex domains of infinite type},
Math. Res. Lett. {\bf 22} (2015), no.~2, 467-484.

\bibitem{hormander:2003}
Lars H{\"o}rmander, The Analysis of Linear Partial Differential Operators~I,
Classics Math. Springer-Verlag, Berlin, 2003.

\bibitem{klimek:1991}
Maciej Klimek, Pluripotential theory, London Mathematical Society Monographs, New Series, {\bf 6},
Oxford Science Publications, The Clarendon Press, Oxford University Press, New York, 1991.

\bibitem{kolodziej:1998}
S{\l}awomir Ko{\l}odziej, {\em The complex Monge--Amp{\`e}re equation}, Acta Math. {\bf 180} (1998),
no.~1, 69-117.

\bibitem{li:2004}
Song-Yin Li, {\em On the existence and regularity of Dirichlet problem
for complex Monge--Amp{\`e}re equations on weakly pseudoconvex domains},
Calc. Var. Partial Differential Equations {\bf 20} (2004),
no.~2, 119-132.

\bibitem{nguyen:2018}
Ngoc Cuong Nguyen, {\em On the H{\"o}lder continuous subsolution problem for the
complex Monge--Amp{\`e}re equation}, Calc. Var. Partial Differential Equations {\bf 57}
(2018), no.~1, Paper No.~8, 15 pp.

\bibitem{nguyen:2020}
Ngoc Cuong Nguyen, {\em On the Hölder continuous subsolution problem for the
complex Monge--Amp{\`e}re equation, II}, Anal. PDE {\bf 13} (2020), no.~2, 435-453.

\bibitem{phong-song-sturm:2012}
D.H. Phong, Jian Song, and Jacob Sturm, {\em Complex Monge--Amp{\`e}re equations}, Surveys
in differential geometry, Vol. XVII, 327-410,
Surv. Differ. Geom., {\bf 17}, Int. Press, Boston, MA, 2012.

\bibitem{sibony:1987}
Nessim Sibony, {\em Une classe de domaines pseudoconvexes}, Duke Math. J.
{\bf 55} (1987), no. 2, 299-319.

\bibitem{walsh:1968}
J. B. Walsh, {\em Continuity of envelopes of plurisubharmonic functions},
J. Math. Mech. {\bf 18} (1968/69), 143-148.

\end{thebibliography}

\documentclass[11pt, reqno]{amsart}

\usepackage{amssymb,latexsym,amsmath,amsfonts,amsthm}
\usepackage{mathrsfs}
\usepackage{enumitem}
\usepackage[usenames]{color}
\usepackage{hyperref}
\usepackage{comment}

\allowdisplaybreaks

\numberwithin{equation}{section}

\definecolor{DPurple}{rgb}{0.46,0.2,0.69}

\theoremstyle{definition}
\newtheorem{definition}{Definition}[section]

\theoremstyle{remark}
\newtheorem{remark}[definition]{Remark}

\theoremstyle{plain}
\newtheorem{theorem}[definition]{Theorem}
\newtheorem{result}[definition]{Result}
\newtheorem{lemma}[definition]{Lemma}
\newtheorem{proposition}[definition]{Proposition}

\newcommand{\zbar}{\overline{z}}

\newcommand{\cHess}{\mathfrak{H}_{\raisebox{-2pt}{$\scriptstyle {\mathbb{C}}$}}}

\newcommand{\dbar}{\overline\partial}
\newcommand{\opCMA}{(dd^c(\bcdot))^n}

\newcommand{\bdy}{\partial}

\newcommand{\cont}{\mathcal{C}(\overline{\Omega})}
\newcommand{\smoo}{\mathcal{C}}

\newcommand{\omgclass}{\mathcal{C}_{\omega}(\overline{\Omega})}
\newcommand{\Bloclass}{\widehat{\mathscr{B}}(\phi, f)}

\newcommand{\bcdot}{\boldsymbol{\cdot}}

\newcommand{\V}{\mathcal{V}}

\newcommand{\Omegabar}{\overline{\Omega}}
\newcommand{\omegatau}{\omega(\|\tau\|)}

\newcommand{\Z}{\mathbb{Z}}
\newcommand{\N}{\mathbb{N}}

\newcommand{\Cn}{\mathbb{C}^n}

\newcommand{\C}{\mathbb{C}} 
\newcommand{\R}{\mathbb{R}}

\begin{document}

\title[H{\"o}lder solutions to the complex Monge--Amp{\`e}re equation]{A necessary and sufficient condition for H{\"o}lder-class solutions to the complex Monge--Amp{\`e}re equation}

\author{Annapurna Banik}
\address{Tata Institute of Fundamental Research,
Centre for Applicable Mathematics, Bengaluru 560065, India}
\email{banikannapurna01@gmail.com, annapurna25@tifrbng.res.in}

\begin{abstract}
We provide a necessary and sufficient condition for the existence of H{\"o}lder continuous
solutions to the complex Monge--Amp{\`e}re equation on bounded domains in $\mathbb{C}^n$.
This condition is motivated by a paper by S.-Y.~Li. We also
prove a result on the regularity of solutions to the complex Monge--Amp{\`e}re equation
on general $B$-regular domains in $\mathbb{C}^n$.
\end{abstract}

\keywords{Complex Monge--Amp{\`e}re equation, H{\"o}lder regular solution, $B$-regular domains.} 
\subjclass[2020]{Primary: 32W20, 32U05 ; Secondary: 32U10, 32T40}

\maketitle

\vspace{-5mm}

\section{Introduction and statement of main results}\label{sec:intro}

Let $\Omega \varsubsetneq \Cn$, $n\geq 2$, be a bounded
domain. Given two functions $\phi \in \smoo(\bdy\Omega; \R)$ and 
$f\in \smoo(\Omega)$, $f\geq 0$, the \emph{Dirichlet problem for the complex
Monge--Amp{\`e}re equation} is the non-linear boundary-value problem that seeks
a function $u \in \smoo(\overline\Omega; \R)$ such that
\begin{align} \label{eqn:D-P_CMA}
  (dd^c{u})^n\,
  :=\,\underbrace{dd^c{u}\wedge\dots \wedge dd^c{u}}_{n \ \text{factors}}\,
  &=\,f\V_n, \ \ u \in {\rm psh}(\Omega) \cap \cont, \\
  u|_{\bdy\Omega}\,&=\,\phi. \notag
\end{align}
Here, $d = (\bdy + \dbar)$ is the exterior derivative, $d^c := i(\bdy - \dbar)$, and
$\V_n$ is defined as
$$
  \V_n := (i/2)^n(dz_1\wedge d\zbar_1)\wedge\dots
  \wedge (dz_n\wedge d\zbar_n).
$$
When $u|_{\Omega}\notin \smoo^2(\Omega; \R)$, the left-hand side of
\eqref{eqn:D-P_CMA} must be interpreted as a current of bidegree $(n, n)$.
That this makes sense when $u\in {\rm psh}(\Omega)\cap \smoo(\Omega)$ was established by
Bedford--Taylor in their seminal paper \cite{bedford-taylor:1976}, where they established
an existence and uniqueness theorem
for \eqref{eqn:D-P_CMA} with continuous data $(f, \phi)$ for strongly pseudoconvex domains.
Later, B\l{}ocki \cite{blocki:1996} showed that such a theorem holds true in a much larger class of
domains, namely, $B$-regular domains (see Section~\ref{subsec:defn-prpn} for the definition and
further discussion). There is extensive literature on the complex Monge--Amp{\`e}re equation
(too numerous to discuss here) in complex analysis, and also in complex geometry;
the reader is referred to \cite{bedford-taylor:1976, caffarelli-kohn-nirenberg-spruck:1985, caffarelli-nirenberg-spruck:1985, kolodziej:1998,
guedj-zeriahi:2017} and the references therein.
\smallskip

The regularity theory for the complex Monge--Amp{\`e}re equation is of great interest.
Loosely speaking, this concerns determining\,---\,assuming $\bdy \Omega$
``nice'' enough\,---\,when the solution to \eqref{eqn:D-P_CMA} has some regularity
(e.g., H{\"o}lder continuity), {\em for all} sufficiently regular data $(f, \phi)$.
The first such result was proved by Bedford--Taylor \cite{bedford-taylor:1976} for $\Omega$
strongly pseudoconvex. We refer the reader to \cite{coman:1997, li:2004,
guedj-kolodziej-zeriahi:2008,
charabati:2015, ha-khanh:2015, baracco-khanh-pinton-zampieri:2016,
nguyen:2018, nguyen:2020} for further work on this subject.
There are many applications of the complex Monge--Amp{\`e}re equation
in complex geometry; see the survey \cite{phong-song-sturm:2012}. There have been
relatively fewer applications of the complex Monge--Amp{\`e}re equation on domains
$\Omega\varsubsetneq \Cn$ outside the realm of PDEs (for $\Omega$ not the unit ball in
$\C^n$). This is largely due to the need for regularity results for solutions to
\eqref{eqn:D-P_CMA}; see \cite{li:2004, banik-bharali:2024, bharali-masanta:2025} for some
applications on domains in $\Cn$. 
\smallskip

An interesting advance in this subject was made by S.-Y.~Li \cite{li:2004}, who proved a
H{\"o}lder-regularity result\,---\,in the spirit of the one by Bedford--Taylor
\cite{bedford-taylor:1976}\,---\,for \eqref{eqn:D-P_CMA} on weakly pseudoconvex domains.
He studied the Dirichlet problem \eqref{eqn:D-P_CMA} on domains that admit a
uniformly strictly plurisubharmonic function $\rho \in {\rm psh}(\Omega) \cap \cont$
(possibly non-smooth), which is the basis for a sufficient condition for the H{\"o}lder regularity
of solutions to \eqref{eqn:D-P_CMA}.
In this paper, we show that such a condition is also necessary.
With these words, we state

\begin{theorem} \label{th:Holder-iff}
Let $\Omega \varsubsetneq \Cn$ be a bounded domain.
The following are equivalent:
\begin{itemize}
    \item[$(a)$] There exist $\varepsilon \in (0,1]$ 
    and a function $\rho \in {\rm psh}(\Omega) \cap \smoo^{0,\,\varepsilon}(\overline{\Omega})$
    such that $ \Omega= {\rho}^{-1}(-\infty, 0)$, $\bdy \Omega = \rho^{-1}\{0\}$,
    and $\rho(\bcdot) - \|\bcdot\|^2$ $ \in {\rm psh}(\Omega)$.

    \smallskip
    
    \item[$(b)$] There exists $\varepsilon \in (0,1]$ such that $\Omega$
    admits a (unique) solution $u$ to the Dirichlet problem
    \eqref{eqn:D-P_CMA} of class $\smoo^{0, \varepsilon}(\overline{\Omega})$,
    for any $\phi \in \smoo^{1, 1}(\bdy \Omega; \R)$
    and for any $f$ such that either $f\equiv 0$ or $f > 0$ with $f \in \smoo^{0,\,\varepsilon}(\overline{\Omega})$.
\end{itemize}
\end{theorem}

The condition $(a)$ is strongly motivated by Li \cite{li:2004} as alluded to above. In fact,
the ``harder'' implication $(a)\!\Rightarrow\!(b)$ in Theorem~\ref{th:Holder-iff} is
{\em essentially} \cite[Corollary~2.2]{li:2004}. However, we shall prove a useful
generalization of \cite[Corollary~2.2]{li:2004}\,---\,see
Theorem~\ref{th:general-reg-CMA}\,---\,from which this implication also follows. An essential
discussion on these matters is deferred to
Remark~\ref{rem:detailed-proof}. Domains that satisfy condition~$(a)$ in
Theorem~\ref{th:Holder-iff} are abundant; this will be discussed in Section~\ref{sec:two-proofs}.%
\smallskip

We must mention that in order to prove this theorem, no assumptions on the
regularity of $\bdy \Omega$ are required a priori. In what follows,
$\phi \in \smoo^{1, 1}(\bdy \Omega)$ means that it is the restriction to $\bdy \Omega$ of a
$\mathcal{C}^{1,1}$-smooth function defined on an open neighbourhood of $\bdy \Omega$.
Also note,
condition~$(a)$ (and, equivalently, condition~$(b)$) implies that $\Omega$ is pseudoconvex.
Since Theorem~\ref{th:Holder-iff} focuses on $f\equiv 0$ or $f > 0$ and
$\phi \in \smoo^{1, 1}(\bdy \Omega)$, two remarks are in order:

\begin{remark} \label{rem:density-data}
We can establish a more general result than Theorem~\ref{th:Holder-iff} by
relaxing the condition on $f$, but doing so results in a statement involving
$f^{1/n}$ that is more technical, while we prefer the more aesthetic result.
Moreover, we are motivated by the applications of the regularity theory for
the complex Monge--Amp{\`e}re equation outside the realm of PDEs, where the
conditions on $f$ as stated in Theorem~\ref{th:Holder-iff} are more frequently
encountered (e.g., $f \equiv 0$ relates to the Perron--Bremermann envelope).
Furthermore, the case when $f \geq 0$ will be discussed in connection with
a more general result as alluded to above\,---\,see Theorem~\ref{th:general-reg-CMA} (and also Theorem~\ref{th:B-reg-CMA}).
\end{remark}

\begin{remark} \label{rem:bdy-value-data}
The choice for the regularity of the boundary data $\phi$ in Theorem~\ref{th:Holder-iff}
is natural. Firstly, one sees from the proofs in \cite{li:2004, guedj-kolodziej-zeriahi:2008}
that the description for a class of boundary data that are strictly less regular
than $\smoo^{1,1}(\bdy \Omega)$, and for which the solution to the Dirichlet problem
\eqref{eqn:D-P_CMA} is in H{\"o}lder class, would be quite technical. Secondly,
in most applications of the regularity theory for the Dirichlet problem outside
the realm of PDEs, one usually works with $\phi \in \smoo^{\infty}(\bdy \Omega)$\,---\,see
\cite{li:2004, phong-song-sturm:2012, banik-bharali:2024, bharali-masanta:2025},
for instance.
\end{remark}

In addition to proving one of the implications of Theorem~\ref{th:Holder-iff},
Theorem~\ref{th:general-reg-CMA} has an amusing consequence, which is our second main
result. Suppose $\Omega \varsubsetneq \Cn$ is a $B$-regular domain. As mentioned earlier in this
section, $\Omega$ admits a unique solution to \eqref{eqn:D-P_CMA} for all
admissible $(f, \phi)$; see Result~\ref{res:B-reg_CMA-soln}. However, in general,
it was not known whether these solutions are H{\"o}lder continuous, even when
$f, \ \phi$ are very regular. Recently, Bharali--Masanta \cite{bharali-masanta:2025}
provided a negative result in this direction. See Section~\ref{sec:two-proofs}
for more details, which will also suggest that it is meaningful to ask:
{\em is there a (concave) modulus of continuity $\omega$\,---\,that depends only on
$\Omega$\,---\,such that \eqref{eqn:D-P_CMA} has a solution in ${\rm psh}(\Omega) \cap \omgclass$
{\bf for all} admissible data $(f, \phi)$?} (See \eqref{eqn:omega-class} for the definition of
$\omgclass$.)
Theorem~\ref{th:general-reg-CMA} addresses this question and leads to an affirmative answer
as stated below. To the best of our knowledge,
no such regularity result for solutions to the complex Monge--Amp{\`e}re equation
on {\em general} $B$-regular domains is known in the literature.
\begin{theorem} \label{th:B-reg-CMA}
Let $\Omega \varsubsetneq \Cn$ be a $B$-regular domain. Then, there exists a modulus of continuity
$\omega$ such that $\Omega$ admits a solution $u$ to \eqref{eqn:D-P_CMA}
of class $\smoo_{\omega}(\overline{\Omega})$ for any $\phi \in \smoo^{1,1}(\bdy \Omega; \R)$
and for any $f \geq 0$ satisfying
$f^{1/n} \in \smoo_{\omega}(\overline{\Omega})$.
\end{theorem}

\section{Analytical preliminaries}\label{sec:premln}

In this section, we note some important definitions and facts that are relevant to
our proofs.%
\smallskip

\subsection{Common notations} \label{subsec:notations}
We write down a list of recurring notations (some of which were used without
comment in the previous section). In this list, $\Omega$ denotes a domain in $\Cn$.
\begin{enumerate}
  \item ${\rm psh}(\Omega)$ denotes the class of all plurisubharmonic functions
  defined on $\Omega$.
  \smallskip
  
  \item Given a $\smoo^2$-smooth function $\psi:\Omega \to \C$, $(\cHess{\psi})(z)$ is the
  complex Hessian of $\psi$ at $z \in \Omega$.%
  \smallskip
  
  \item For $v \in \R^d$, $\|v\|$ denotes the Euclidean norm. Given a set non-empty set
  $S \subset \R^d$, ${\rm diam}(S): = \sup \{\|a-b\|: a, b \in S\}$.
  \smallskip
  
  \item Let $S (\neq \emptyset) \subset \Cn$. For fixed $j \in \N$, $\beta \in (0,1]$,
  $\smoo^{j,\,\beta}(S)$ is the class of all functions that are continuously
  differentiable to order $j$ and whose $j$-th partial derivatives satisfy a uniform H{\"o}lder
  condition with exponent $\beta$. $\smoo(S)$ denotes the class of continuous functions.
  The subclasses $\smoo^{j,\,\beta}(S; \R) \subset \smoo^{j,\,\beta}(S)$ and
  $\smoo(S; \R) \subset \smoo(S)$ consist of $\R$-valued functions.
\end{enumerate}

\subsection{The complex Monge--Amp{\`e}re operator}
We briefly recall some basic facts about the complex Monge--Amp{\`e}re operator
$\opCMA$, as featured in \eqref{eqn:D-P_CMA}, which will be relevant to the proof of
Theorem~\ref{th:general-reg-CMA}. Let $\Omega \subset \Cn$ be a domain.
Let $\mathcal{M}_{n}(\Omega)$ denotes the class of currents on $\Omega$
of bidegree $(n, n)$ and of order $0$, equipped with  the usual topology
of weak convergence. It is well known that elements of $\mathcal{M}_{n}(\Omega)$
are identified with $(n,n)$-differential forms with measure coefficients.
\smallskip

Firstly, an easy computation yields, for any
$\smoo^2$-smooth function $U$ on $\Omega$,
\begin{equation} \label{eqn:An}
  (dd^c{U})^n\,
  :=\,\underbrace{dd^c{U}\wedge\dots \wedge dd^c{U}}_{n \ \text{factors}}
  = A_n (\det \cHess U ) \V_n, \quad \text{where } A_n:= 4^n n \, !\, .
\end{equation}
Suppose $U \in {\rm psh}(\Omega) \cap \smoo^2(\Omega)$, then $\det \cHess U \geq 0$. Therefore,
by \eqref{eqn:An},
$ (dd^c{U})^n$ is a positive $(n,n)$-current and hence is in $\mathcal{M}_{n}(\Omega)$.
Bedford--Taylor \cite[Proposition~2.3]{bedford-taylor:1976} showed that there exists
a unique continuous operator
$$
  (dd^c(\bcdot))^n: {\rm psh}(\Omega) \cap \smoo(\Omega) \to \mathcal{M}_{n}(\Omega)
$$
that extends $(dd^c(\bcdot))^n|_{{\rm psh}(\Omega) \cap \smoo^2(\Omega)}$ (which is
to be understood in the usual sense as in \eqref{eqn:An}).
As ${\rm psh}(\Omega) \cap \smoo^2(\Omega)$ is dense in ${\rm psh}(\Omega) \cap \smoo(\Omega)$, $(dd^c{u})^n$ is a positive
$(n,n)$-current for every $u \in {\rm psh}(\Omega) \cap \smoo(\Omega)$.
For more details about $\opCMA$, see, for instance, \cite{bedford-taylor:1976},
\cite{blocki:1996}, \cite{guedj-zeriahi:2017}.
\smallskip

\subsection{The complex Monge--Amp{\`e}re equation on $B$-regular domains}
\label{subsec:defn-prpn} 
We begin this section with a result of Sibony \cite{sibony:1987} that encodes the
definition of $B$-regular domains.
\begin{result}[paraphrasing {\cite[Th{\'e}or{\`e}me~2.1]{sibony:1987}}]
\label{res:B-reg_equiv-prop}
Let $\Omega \varsubsetneq \Cn$ be a bounded domain. The following are equivalent.
\begin{itemize}
  \item[$(a)$] For each $\zeta \in \bdy \Omega$, $\Omega$ admits a strong plurisubharmonic barrier,
  i.e., there exists $v_{\zeta} \in {\rm psh}(\Omega)$ such that 
  $$
    \lim_{\Omega \ni z \to \zeta} v_{\zeta}(z)=0 \text{   and   }
    \limsup v_{\zeta}|_{\Omegabar \setminus \{\zeta\} } < 0.
  $$
  
  \item[$(b)$] $\Omega$ admits an exhaustion $\psi \in {\rm psh}(\Omega) \cap \smoo^{\infty}(\Omega) \cap \cont$
  with $\psi|_{\bdy \Omega}\equiv 0$ such that
  $$
    \sum_{j,k =1}^{n} \frac{\bdy^2 \psi}{\bdy z_j\bdy \overline{z}_k} v_j \overline{v}_k \geq \|v\|^2,
    \quad \forall v=(v_1, \cdots, v_n) \in \Cn.
  $$

  \item[$(c)$] For every function $f \in \smoo(\Omegabar; \R)$, there exists $v \in
  {\rm psh} (\Omega) \cap \cont$ such that $v|_{\bdy \Omega} = f$.
\end{itemize}
\end{result}

\begin{definition}\label{defn:B-reg}
A bounded domain $\Omega \varsubsetneq \Cn$ is said to be \emph{$B$-regular} if it satisfies
any of the equivalent conditions stated in Result~\ref{res:B-reg_equiv-prop}.
\end{definition}

\begin{remark}
In \cite{sibony:1987}, Sibony introduced the notion of {\em $B$-regularity}
of the boundary $\bdy \Omega$ of a bounded domain $\Omega \varsubsetneq \Cn$.
In fact, $\bdy \Omega$ being $B$-regular, as defined in \cite{sibony:1987},
is equivalent to each of the conditions stated in
Result~\ref{res:B-reg_equiv-prop} under the assumption that $\bdy \Omega$ is
$\smoo^1$-smooth\,---\,see \cite[Th{\'e}or{\`e}me~2.1]{sibony:1987}. However, we retain
Definition~\ref{defn:B-reg} in its form as it will be useful for our proofs. In particular,
some of our arguments rely on the results in \cite{blocki:1996}, where $B$-regularity
of $\Omega$ is defined as in Definition~\ref{defn:B-reg}\,---\,see \cite[p.~728]{blocki:1996}.
\end{remark}
\smallskip

We shall now state the result by B\l{}ocki \cite{blocki:1996}, as alluded to in
Section~\ref{sec:intro}, regarding the existence and uniqueness of solution to \eqref{eqn:D-P_CMA}
on $B$-regular domains. To do so, we first define the following Perron--Bremermann family:
given $\phi \in \smoo(\bdy \Omega; \R)$, $f \in \smoo(\overline{\Omega})$ with $f \geq 0$,
\begin{equation}\label{eqn:Blo-PB-envp}
\widehat{\mathscr{B}}(\phi, f) := \{ v \in {\rm psh}(\Omega) \cap \smoo(\overline{\Omega}):
v|_{\bdy \Omega}= \phi, \ (dd^c v)^n \geq f \V_n \}.
\end{equation}

\begin{result}[paraphrasing {\cite[Theorem~4.1]{blocki:1996}}] \label{res:B-reg_CMA-soln}
Let $\Omega \varsubsetneq \Cn$ be a $B$-regular domain. Let $\phi \in \smoo(\bdy \Omega; \R)$ and
$f \in \smoo(\overline{\Omega})$ with $ f \geq 0$. Then, there exists a unique solution
$u \in {\rm psh}(\Omega) \cap \smoo (\overline{\Omega})$ to the Dirichlet problem
\eqref{eqn:D-P_CMA}, where the solution can be taken of the form
$$
  u(z) = \sup \{ v(z): v \in \widehat{\mathscr{B}} (\phi, f) \}.
$$
\end{result}

The above result is strongly related to \cite[Theorem~8.3]{bedford-taylor:1976}.
We mention that certain other Perron--Bremermann families corresponding to continuous data
$(f, \phi)$\,---\,which are larger than the one defined by
\eqref{eqn:Blo-PB-envp}\,---\,are also considered in the literature to
describe the solution to \eqref{eqn:D-P_CMA}; see, for example,
\cite{bedford-taylor:1976, li:2004, ha-khanh:2015}. We note a couple of such families
that appear frequently:
\begin{align} \label{eqn:PB-envp}
  \mathscr{F}(\phi, f) &:= \{ v \in {\rm psh}(\Omega) \cap \smoo(\overline{\Omega}): v|_{\bdy \Omega} \leq \phi,
  \ (dd^c v)^n \geq f \V_n \},  \notag \\
  \mathscr{B}(\phi,f ) &:= \{ v \in {\rm psh}(\Omega) \cap \smoo(\Omega): \limsup v|_{\bdy \Omega} \leq \phi,
  \ (dd^c v)^n \geq f \V_n \}.
\end{align}
The following proposition asserts that the upper envelopes of all Perron--Bremermann
families defined above coincide in the case of $B$-regular domains. We conclude this
section stating the result; the proof is straightforward and will be omitted.
\begin{proposition} \label{prpn:eq-envelopes}
Let $\Omega \varsubsetneq \Cn$ be a $B$-regular domain. Let $\phi \in \smoo(\bdy \Omega; \R)$
and $f \in \smoo(\overline{\Omega})$ with $ f \geq 0$. Then,
$$
  \sup \{ v(z): v \in \Bloclass \} = \sup \{ v(z): v \in \mathscr{F}(\phi, f) \}
  = \sup \{ v(z): v \in \mathscr{B}(\phi, f) \} \quad \forall z \in \Omega.
$$   
\end{proposition}

\section{The proof of Theorem~\ref{th:general-reg-CMA}}\label{sec:general-reg-result}   
\smallskip

This section is devoted to proving Theorem~\ref{th:general-reg-CMA}, the more general as
alluded to in Section~\ref{sec:intro}. Among other things, this result provides
the proof of one of the implications in Theorem~\ref{th:Holder-iff}.
We need the following definitions before we state this result.

\begin{definition}\label{defn:mod-cont}
A function $\omega: [0,\infty)\to [0,\infty)$ is called a \emph{modulus of
continuity} if it is concave, monotone increasing, and such that
$\lim_{x\to 0^+}\omega(x) = \omega(0) = 0$.
\end{definition}
With $\omega$ as above, we define
\begin{equation}\label{eqn:omega-class}
  \smoo_{\omega}(\overline{\Omega}):= \{\psi\in \smoo(\overline{\Omega}): \exists C>0 \text{ such that }
  |\psi(x)- \psi (y)| \leq C \omega (\|x-y \|) \ \forall x, y \in \overline{\Omega}\}.
\end{equation}

\begin{theorem}\label{th:general-reg-CMA}
Let $\Omega \varsubsetneq \Cn$ be a bounded domain. Assume that there exists a modulus of
continuity $\omega$ and a function $\rho \in {\rm psh}(\Omega) \cap
\smoo_{\omega}(\overline{\Omega}) $ such that $ \Omega= {\rho}^{-1}(-\infty, 0)$,
$\bdy \Omega = \rho^{-1}\{0\}$, and $\rho(\bcdot) -  \|\bcdot\|^2 \in {\rm psh}(\Omega)$.
Then, $\Omega$ admits a solution $u$ to \eqref{eqn:D-P_CMA} of class
$\smoo_{\omega}(\overline{\Omega})$ for any $\phi \in \smoo^{1,1}(\bdy \Omega; \R)$ and
for any $f \geq 0$ satisfying $f^{1/n} \in \smoo_{\omega}(\overline{\Omega})$.
\end{theorem}

\begin{remark} \label{rem:detailed-proof}
As discussed in Section~\ref{sec:intro}, the above theorem is strongly motivated by the work
of Li \cite{li:2004}. In particular, our
sufficient condition involving a function
$\rho \in {\rm psh}(\Omega) \cap \cont$ is inspired by \cite[Corollary~2.2]{li:2004}.
A careful reading of the proofs in \cite{li:2004} indicates that the proof of
\cite[Corollary~2.2]{li:2004} relies on \cite[Proposition~2.1]{li:2004}.
Despite the connection of Theorem~\ref{th:general-reg-CMA} with \cite{li:2004}
and a {\em seeming} similarity in the proofs, we feel that Theorem~\ref{th:general-reg-CMA}
requires a detailed proof. This is because there seems to be a slight lack of clarity in the proof
of \cite[Proposition~2.1]{li:2004}. For instance:
\begin{enumerate}
  \item[(a)] What is referred to as ``$v$'' in the statement of \cite[Proposition~2.1]{li:2004}
  is not the function $v$ defined at the beginning of the proof of this proposition; 
  it is\,---\,per the discussion on \cite[p.~122]{li:2004}\,---\,the upper-semicontinuous regularization
  of a certain function ``$u$'' defined on p.~122. However, as defined, this $u$ is manifestly
  H{\"o}lder continuous and needs no regularization. For the same reason, furthermore, the subsequent
  long argument for why $u$ is H{\"o}lder continuous on $\Omegabar$ is unnecessary.
  \smallskip

  \item[(b)] The above-mentioned function $u$ is given as the upper envelope of a certain family; see
  \cite[p.~122]{li:2004}. It is not clear why
  $(dd^c u)^n = f \V_n$, despite the reference to \cite{bedford-taylor:1976}, because the latter
  family seems too small for the intended purpose. 
\end{enumerate}
\end{remark}

Adapting the ideas of \cite[Proposition~2.1]{li:2004} and
\cite[Theorem~6.2]{bedford-taylor:1976}, we give a thorough proof of
Theorem~\ref{th:general-reg-CMA}. The next two lemmas will be useful for the proof.

\begin{lemma}\label{l:aux-psh}
Let $\Omega \varsubsetneq \Cn$ be a bounded domain and let $\psi \in \smoo^{1,1}(\overline{\Omega}; \R)$.
Suppose that $\Omega$ admits $\rho \in {\rm psh}(\Omega) \cap \cont$ such that
$ \Omega= {\rho}^{-1}(-\infty, 0)$,
$\bdy \Omega = \rho^{-1}\{0\}$, and $\rho(\bcdot) -  \|\bcdot\|^2 \in {\rm psh}(\Omega)$.
Then, there exists $K > 0$, large enough, such that $\psi (\bcdot) + K \rho (\bcdot ) \in 
{\rm psh} (\Omega) \cap \smoo (\overline{\Omega})$.
\end{lemma}

\begin{lemma}[Ha--Khanh, {\cite[Lemma~3.4]{ha-khanh:2015}}] \label{l:positive-matrix}
Let ${\sf M} \in M_{n \times n}(\C)$ be a non-negative, $n \times n$ Hermitian matrix
and let $\beta \geq 0$. Then,
$$
  \det ( {\sf M} + \beta \, {\sf I}_n ) \geq \sum_{k=0}^n {\beta}^k (\det {\sf M})^{\frac{n-k}{n}},
$$
where ${\sf I}_n \in M_{n \times n}(\C)$ denotes the identity matrix.
    
\end{lemma}

\begin{proof}[The proof of Theorem~\ref{th:general-reg-CMA}]
We fix a $\phi \in \smoo^{1,1}(\bdy \Omega; \R)$ and an $f \geq 0$ satisfying
$f^{1/n} \in \smoo_{\omega}(\overline{\Omega})$. As we recall from Section~\ref{subsec:defn-prpn},
\begin{equation} \label{eqn:Blo-env}
  \widehat{\mathscr{B}}(\phi, f) = \{ v \in {\rm psh}(\Omega) \cap \smoo(\overline{\Omega}):
  v|_{\bdy \Omega}= \phi, \ (dd^c v)^n \geq f \V_n \}.
\end{equation}
Let $u$ be the upper envelope of this Perron--Bremermann family, i.e.,
\begin{align} \label{eqn:soln-candidate}
  u(z) &= \sup \{ v(z): v \in \widehat{\mathscr{B}} (\phi, f) \}, \ z \in \overline{\Omega}.
\end{align}
We shall present the proof in the following four steps. In the first step, we explain why the
function $u$ as defined in \eqref{eqn:soln-candidate} solves \eqref{eqn:D-P_CMA}. In the
subsequent steps, we will show that $u \in \omgclass$.%
\medskip

\noindent{\textbf{Step 1.}} \emph{Showing that $u$ solves the Dirichlet problem \eqref{eqn:D-P_CMA} }
\smallskip

\noindent{First,} we note that for each $\zeta \in \bdy \Omega$, $\Omega$ admits a
strong plurisubharmonic barrier $v_{\zeta}$. Define
$$
  v_{\zeta}(z):= - \|z - \zeta\|^2 + \rho(z), \quad \forall z \in \Omegabar.
$$
Clearly, given the assumptions on $\rho$, $v_{\zeta}$ is a strong plurisubharmonic barrier
for each $\zeta \in \bdy \Omega$. So, by Result~\ref{res:B-reg_equiv-prop},
$\Omega$ is a $B$-regular domain. Thus, by Result~\ref{res:B-reg_CMA-soln}, the function
$u \in {\rm psh}(\Omega) \cap \smoo(\overline{\Omega})$ as defined by \eqref{eqn:soln-candidate}
solves the Dirichlet problem \eqref{eqn:D-P_CMA} with data $(f, \phi)$ as fixed above.
\medskip

\noindent{\textbf{Step 2.}} \emph{Showing that there exists $ K >0$ such that
$|u(z) - u (\zeta)| \leq K \omega ( \|z - \zeta\|) $ for all
$(z, \zeta) \in \overline{\Omega} \times \bdy \Omega$}
\smallskip

\noindent{We} extend $\phi$ to a function in $\smoo^{1,1}(\overline{\Omega}; \R)$ and call
it $\phi$ as well. By Lemma~\ref{l:aux-psh}, there is $K>0$ such that both of the functions
$$
  v(z):= \phi(z) + K \rho (z) \quad \text{and} \quad \widetilde{v}(z):= - \phi(z) + K \rho (z)
$$
are in ${\rm psh}(\Omega) \cap \cont$. And moreover, we can choose $K$ so large that
$(dd^cv)^n \geq f \V_n$. Clearly, $v \in \Bloclass$ and hence $v \leq u$. Also,
$u + \widetilde{v} \in {\rm psh}(\Omega) \cap \smoo(\overline{\Omega})$ and
$(u + \widetilde{v})|_{\bdy \Omega}=0$. Therefore,
\begin{equation} \label{eqn:sub-sup-estimates}
  v \leq u \leq - \widetilde{v} \ \text{ on } \overline{\Omega} \quad \text{and} \quad
  u|_{\bdy \Omega}=v|_{\bdy \Omega}= -\widetilde{v}|_{\bdy \Omega} = \phi,
\end{equation}
where the second inequality holds because of the maximum principle for plurisubharmonic functions.
Now, as $\omega$ is concave and $\omega(0)= 0$, there exists $c>0$ such that 
$$
  \omega(x) \geq c x \quad \text{near } 0 \in \R.
$$
So, we must have $\phi \in \omgclass$ and hence $v, \widetilde{v} \in \omgclass$.
Thus, using \eqref{eqn:sub-sup-estimates}, there exists $K_1>0$ such that
\begin{equation} \label{eqn:soln-aux-estimate}
  |u(z) - u (\zeta)| \leq K_1 \omega ( \|z - \zeta\|) \quad
  \forall (z, \zeta) \in \overline{\Omega} \times \bdy \Omega.
\end{equation}

\medskip

\noindent{\textbf{Step 3.}} \emph{Constructing auxiliary functions $V_{\tau}: \Omegabar \to \R$
for $\tau \in \Cn, \ \|\tau\|$ small, such that $V_{\tau} \in \Bloclass$}
\smallskip

\noindent{For} $\tau \in \Cn, \ \|\tau\|$ small enough, we define
\begin{equation} \label{eqn:vtau}
  v_{\tau}(z):= u (z + \tau) + C_f \omegatau \|z\|^2 - K_1 \omegatau - K' \omegatau \quad
  \forall z \in \overline{\Omega - \tau},
\end{equation}
where $C_f>0$ is a constant (that depends only on $f$) to be determined later
(see \eqref{eqn:C_f}), $K_1>0$ is given by $\eqref{eqn:soln-aux-estimate}$, and $K'>0$ is so chosen
that $K' > C_f \sup_{z \in \Omegabar} \|z\|^2$.
\smallskip

Now, using \eqref{eqn:soln-aux-estimate} and \eqref{eqn:vtau}, we get
\begin{align}\label{eqn:vtau-estm}
  v_{\tau}(z) \leq u(z) + (C_f \|z\|^2 - K')\omegatau \leq u(z) \quad
  \forall z \in \bdy (\Omega - \tau) \cap \Omegabar,
\end{align}
where the second inequality is due to the choice of $K'$ as above.
\smallskip

We now appeal to the following standard construction\,---\,which is due to Walsh
\cite{walsh:1968}\,---\,to complete the proof. Define for $\tau \in \Cn, \ \|\tau\|$ small enough,
$$
  V_{\tau}: \Omegabar \to \R, \  
  V_{\tau}(z):=
  \begin{cases}
   u(z) \quad  &\text{if } z \in \Omegabar \setminus (\Omega - \tau), \\
   \max \{u(z), v_{\tau}(z)\} \quad &\text{if } z \in \Omegabar \cap \overline{\Omega - \tau}.
  \end{cases}
$$
In view of \eqref{eqn:vtau-estm}, it is easy to see that $V_{\tau}$ is well-defined
and is in ${\rm psh}(\Omega) \cap \cont$. We claim that $V_{\tau} \in \Bloclass$.
\smallskip

Note, by \eqref{eqn:soln-aux-estimate}, we get $v_{\tau} (\zeta) \leq u (\zeta)$ for each
$\zeta \in \bdy \Omega \cap (\Omega - \tau)$. Therefore,
$$
  V_{\tau} (\zeta) = u(\zeta) = \phi(\zeta) \quad \forall \zeta \in \bdy \Omega,
$$
by the definition of $V_{\tau}$. Thus, to establish the above claim, it remains to show that
$(dd^c V_{\tau})^n \geq f \V_n$. In view of \cite[Proposition~2.8]{bedford-taylor:1976},
to establish the latter, it suffices to show that
\begin{equation} \label{eqn:vtau_MA-estm-final}
  (dd^c v_{\tau})^n \geq f \V_n \quad \text{on } \Omega \cap (\Omega - \tau) .
\end{equation}

We shall now choose $C_f>0$ as introduced in \eqref{eqn:vtau}. Let $c_f >0$ be such that 
\begin{equation} \label{eqn:density-estm}
  |f^{1/n}(x) - f^{1/n}(y)| \leq c_f \, \omega (\|x-y\|) \quad \forall x, \, y \in \Omegabar.
\end{equation}
Let $A_n>0$ be the constant as given by \eqref{eqn:An} so that
$(dd^c U)^n = A_n (\det \cHess{U}) \V_n$ for any $U \in \smoo^2(\Omega)$. Let
\begin{equation} \label{eqn:C_f}
  C_f > \max_{k = 1, \cdots, n} {n \choose k}^{1/k} {A_n}^{-1/n} c_f
\end{equation}
This choice of $C_f$ will ensure that \eqref{eqn:vtau_MA-estm-final} can be established,
as we shall see below in this step.
\smallskip

Let $u_{\epsilon} ( \bcdot), \ \epsilon>0$, be the standard regularization of
$u(\bcdot)$ (see, for instance, \cite[Chapter~2]{klimek:1991}).
Let us take a sequence $\{\epsilon_{\nu}\}$ that decreases to $0$ and write
$u_{\nu}:= u_{\epsilon_{\nu}}$. Technically speaking, in order to deduce
\eqref{eqn:vtau-MA-estm}, we must now fix an arbitrary subsequence of $\{u_{\nu}\}$ but,
without loss of generality (and for simplicity of notation), we shall refer to the latter
as $\{u_{\nu}\}$.
It is a standard argument following from the definitions (see, \cite[Chapters~III-IV]{hormander:2003},
for instance) that there exists a sequence
$\{\nu_j\}\subset \Z_+$ such that the right-hand side of the following equation (where the
limit is stated is in the weak sense) is convergent a.e.
\begin{equation}\label{eqn:vtau-MA-estm-raw1}
  (dd^c v_{\tau})^n = \lim_{j \to \infty} \big(dd^c \big(u_{\nu_j}(\bcdot + \tau)
  + C_f\, \omegatau \|\bcdot\|^2 \big) \big)^n.
\end{equation}
Now, owing to convergence a.e. we have:
\begin{align}
  \lim_{j \to \infty}& \big(dd^c \big(u_{\nu_j}(\bcdot + \tau)
  + C_f\, \omegatau \|\bcdot\|^2 \big) \big)^n \notag \\
  &= A_n \lim_{j \to \infty} \det \big( \cHess{u_{\nu_j}(\bcdot + \tau)}
  + C_f \, \omegatau \, {\sf I}_n \big) \notag \\
  &\geq A_n \lim_{j \to \infty} \Big(\sum_{k=0}^n
  C^k_f \, \omegatau^k \det \big( \cHess{u_{\nu_j}(\bcdot + \tau)}\big)^{\frac{n-k}{n}} \Big) \ \
  \text{(by Lemma~\ref{l:positive-matrix})} \notag \\
  &= f(\bcdot + \tau) + A_n \sum_{k=1}^n A^{\frac{k}{n}-1}_n C^k_f \, \omegatau^k
  f^{\frac{n-k}{n}}(\bcdot + \tau),
  \label{eqn:vtau-MA-estm-raw2}
\end{align}
where the first equality and the first inequality hold a.e. while (due to the dominated
convergence theorem) in \eqref{eqn:vtau-MA-estm-raw2}, the limits are taken
in the weak sense. From \eqref{eqn:vtau-MA-estm-raw1} and \eqref{eqn:vtau-MA-estm-raw2} and the
fact that $\{\epsilon_{\nu}\}$ was arbitrary, we have
\begin{equation}\label{eqn:vtau-MA-estm}
  (dd^c v_{\tau})^n \geq f(\bcdot + \tau) + A_n \sum_{k=1}^n A^{\frac{k}{n}-1}_n C^k_f \, \omegatau^k
  f^{\frac{n-k}{n}}(\bcdot + \tau)
\end{equation}  
(also see \cite[inequality (3.10)]{ha-khanh:2015}).    
Again, by \eqref{eqn:density-estm} and \eqref{eqn:C_f} we have
\begin{align}\label{eqn:density-estm-final}
  f(z) &\leq f(z + \tau) + \sum_{k =1}^{n} {n \choose k} c^k_f \, \omegatau^k \, f^{\frac{n-k}{n}}(z + \tau)
  \notag \\
  &\leq f(z + \tau) + \sum_{k =1}^{n} A^{\frac{k}{n}}_n C^k_f \, \omegatau^k \,
  f^{\frac{n-k}{n}}(z + \tau) \quad  \forall z \in \Omega \cap (\Omega - \tau).
\end{align}
Combining \eqref{eqn:density-estm-final} and \eqref{eqn:vtau-MA-estm}, we can establish
\eqref{eqn:vtau_MA-estm-final}. This implies that $(dd^c V_{\tau})^n \geq f \V_n$ and hence
$V_{\tau} \in \Bloclass$.
\medskip

\noindent{\textbf{Step 4.}} \emph{Showing that $u \in \omgclass$}
\smallskip

\noindent{Let} $z, z' \in \Omegabar$, $\|z-z'\|$ be small enough, and write $\tau:= z'-z$.
Now, $V_{\tau} \in \Bloclass$ by Step~3. Note that $z \in \Omegabar \cap \overline{\Omega - \tau}$.
So we must get 
$$ 
  v_{\tau}(z) \leq V_{\tau}(z)\leq u(z) .
$$
Then, in view of $\eqref{eqn:vtau}$, the above inequality gives
\begin{align*}
  u(z') - u(z) &\leq (K_1+K' )\omegatau  - C_f \omegatau \|z\|^2 \leq (K_1+K' )\, \omega (\|z-z'\|).
\end{align*}
Reversing $z$ and $z'$ and proceeding exactly as above, we conclude that $u \in \omgclass$.
This completes the proof.
\end{proof}

\section{The proofs of Theorem~\ref{th:Holder-iff} and Theorem~\ref{th:B-reg-CMA}}
\label{sec:two-proofs}

Before we prove Theorem~\ref{th:Holder-iff}, we must mention that domains that satisfy condition~$(a)$ in this theorem are abundant. Clearly, any bounded
strongly pseudoconvex domain will satisfy this condition. Moreover, any
smoothly bounded pseudoconvex domain of finite type in $\C^2$, or smoothly bounded convex
domain of finite type in $\C^n$, satisfies condition~$(a)$. This follows from
\cite[Theorem~A]{fornaess-sibony:1989} or \cite[Theorem~4.1]{fornaess-sibony:1989}, as
appropriate, combined with a construction presented in the \emph{proof}
of \cite[Proposition~5.1]{fornaess-sibony:1989}.
(Also see \cite[Remark~1]{li:2004}.) The form that the consequences of
\cite[Theorem~A, Theorem~4.1]{fornaess-sibony:1989} take allows one to add to the list
of domains in $\Cn$, $n\geq 3$, that satisfy condition~$(a)$. The last point will be discussed in a
forthcoming work.

\begin{proof}[The proof of Theorem~\ref{th:Holder-iff}]
The implication $(a) \Rightarrow (b)$ follows from Theorem~\ref{th:general-reg-CMA}. To see this,
take $\omega (t):= t^{\varepsilon}, \ t \geq 0$, and use the fact that
whenever $f \equiv  0$ or $f>0$,
$$
  f \in \smoo^{0,\,\varepsilon}(\overline{\Omega}) \Leftrightarrow
  f^{1/n} \in \smoo^{0,\,\varepsilon} (\overline{\Omega}).
$$

Conversely, assume that $(b)$ holds. Define
$\phi: \bdy \Omega \to \R$ as
$$
  \phi(z):= - \|z\|^2.
$$
Clearly $\phi \in \smoo^{1, 1}(\bdy \Omega)$. Then, by $(b)$, there exist
$\varepsilon \in (0,1]$, and $u_{\phi} \in {\rm psh}(\Omega)
\cap \smoo^{0, \varepsilon}(\overline{\Omega})$ that solves the Dirichlet problem
\eqref{eqn:D-P_CMA} with $\phi$ as defined above, and with $f\equiv 0$ on
$\overline{\Omega}$. Now, define
\begin{equation} \label{eqn:psh-candidate}
  \rho (z): = u_{\phi}(z) +  \|z\|^2 \quad \forall z \in \overline{\Omega}.   
\end{equation}

Clearly, $\rho \in {\rm psh}(\Omega) \cap \smoo^{0, \varepsilon}(\overline{\Omega})$
and $ \rho(\bcdot) - \|\bcdot\|^2  \in {\rm psh}(\Omega)$. Also, it follows from
\eqref{eqn:psh-candidate} that $\rho|_{\bdy \Omega}\equiv 0$. By the maximum principle,
$\rho < 0$ on $\Omega$. This proves $(a)$.
\end{proof}

Finally, we deduce the proof of Theorem~\ref{th:B-reg-CMA} from
Theorem~\ref{th:general-reg-CMA}. Before doing so, we discuss
whether H{\"o}lder regularity of solutions to \eqref{eqn:D-P_CMA} can be expected
for all admissible data on general $B$-regular domains. The answer is in the \textbf{negative}, as
shown by Bharali--Masanta
\cite{{bharali-masanta:2025}}. They constructed a class of $B$-regular domains
for which \eqref{eqn:D-P_CMA} cannot admit H{\"o}lder regular solutions
for arbitrary data as prescribed by \eqref{eqn:D-P_CMA}, even for $\phi$ that is highly
regular. But their arguments suggest even more.
The following statement is a combination of \cite[Theorem~1.1]{bharali-masanta:2025} and
\cite[Example~2.4]{bharali-masanta:2025}.

\begin{result}[Bharali--Masanta, \cite{bharali-masanta:2025}] \label{res:non-Holder_CMA}
There exist $B$-regular domains $\Omega \varsubsetneq \Cn, n \geq 2$, with $\smoo^2$-smooth
boundary having the following property: there exist functions $\phi: \bdy \Omega \to \R$ that are
restrictions of $\smoo^{\infty}$-smooth functions defined on neighbourhoods of $\bdy \Omega$ such that,
for any $f \in \smoo(\Omegabar), \, f \geq 0$, the unique solution to the Dirichlet problem
\eqref{eqn:D-P_CMA} does not belong to $\smoo^{0, \varepsilon}(\Omegabar)$
for any $\varepsilon \in (0, 1]$.  
\end{result}

In fact, the \emph{proof} of \cite[Theorem~1.1]{bharali-masanta:2025} reveals that,
given any concave modulus of continuity $\widetilde{\omega}$, there exists a class of $B$-regular
domains with $\smoo^2$-smooth boundary
for which \eqref{eqn:D-P_CMA} cannot admit solutions in $\smoo_{\widetilde{\omega}}(\Omegabar)$
for arbitrary data $(f,\phi)$, even for $\phi$ highly regular. It is this that motivates the question
raised in Section~\ref{sec:intro}.
Theorem~\ref{th:B-reg-CMA} can be viewed as a counterpoint to the above discussion.

\begin{proof}[The proof of Theorem~\ref{th:B-reg-CMA}]
As $\Omega$ is $B$-regular, by Result~\ref{res:B-reg_equiv-prop},
it admits an exhaustion $\rho \in {\rm psh}(\Omega) \cap \smoo^{\infty}(\Omega)
\cap \smoo(\overline{\Omega})$ with $\bdy \Omega = \rho^{-1}\{0\}$ such that
$$
  \sum_{j,k =1}^{n} \frac{\bdy^2 \rho}{\bdy z_j\bdy \overline{z}_k} v_j \overline{v}_k \geq \|v\|^2,
  \quad \forall v=(v_1, \cdots, v_n) \in \Cn.
$$
Clearly, $\rho(\bcdot) - \|\bcdot\|^2 \in {\rm psh}(\Omega)$. Also,
by the maximum principle $ \Omega= {\rho}^{-1}(-\infty, 0)$.
\smallskip

Now, define $\omega_{\rho}:[0, \infty) \to [0, \infty) $ as
$$
  \omega_{\rho} (r):=
  \begin{cases}
      \sup \big\{|\rho(x) - \rho(y)| : x, y \in \overline{\Omega}, \ \|x-y\| \leq r \big\}, 
      \smallskip
      
      &\text{if } 0 \leq r \leq  {\rm diam}(\overline{\Omega}), \\
      \sup
      \big\{|\rho(x) - \rho (y)|: {x, y \in \overline{\Omega}} \big\},
      &\text{if } r \geq {\rm diam}(\overline{\Omega}).
  \end{cases}
$$
Let $\omega$ be the least concave majorant of $\omega_{\rho}$\,---\,see
\cite[Chapter~2, Section~6]{DeVore-Lorentz:1993} for details. It follows from
\cite[Lemma~6.1]{DeVore-Lorentz:1993} that $\omega$ is a (concave) modulus of
continuity and $\rho \in \smoo_{\omega}(\overline{\Omega})$.
The result now follows immediately from Theorem~\ref{th:general-reg-CMA}.
\end{proof}

\section*{Acknowledgements}
I am grateful to Prof. Gautam Bharali for introducing me to this subject and for suggesting one of
the problems in this paper. I also thank him for many valuable discussions and for his help in improving the
exposition of this article. This work is supported by a postdoctoral fellowship from Tata Institute
of Fundamental Research, Centre for Applicable Mathematics.

\end{document}